# Homophily Within and Across Groups

### v2

@@ IV Percolation
IV.1 Colored Percolation
Figure 3: Percolation Phase Transition in Networks with Different within- and across-homophily Values. Local mixing patterns affect the size of the giant component and the percolation threshold. Panels (a–d) show the percolation transition in the size of the giant component for networks composed of 2- and 6-cliques ($G=G_{2}\cup G_{6}$). Each curve corresponds to a different pair of homophily levels, $h_{2}$ and $h_{6}$, and tracks the growth of the giant component as the red–red link activation probability $\pi_{\mathrm{rr}}$ increases. Link activation settings vary across panels: (a) $\pi_{\mathrm{rb}}=\pi_{\mathrm{bb}}=0.2$; (b) $\pi_{\mathrm{rb}}=0$, $\pi_{\mathrm{bb}}=0.2$; (c) $\pi_{\mathrm{rb}}=0$, $\pi_{\mathrm{bb}}=0$; and (d) $\pi_{\mathrm{rb}}=0.2$, $\pi_{\mathrm{bb}}=0$. All networks are generated with $N=500{,}000$ nodes, average degree six, red node fraction $n_{\mathrm{r}}=0.6$, link composition $\alpha=0.5$, and Coleman index $h=0.3$. Results are averaged over 20 realizations. In (a), all networks remain in the supercritical regime regardless of $\pi_{\mathrm{rr}}$, indicating that an average degree of six is sufficient for global connectivity under moderate homophily. By contrast, in (b) and (d), the percolation order of the curves reverses: the configuration that percolates last in (b) is the first to do so in (d). This reversal highlights how activating within- versus across-group links reshapes the percolation threshold, even when the overall homophily index remains fixed.
In a standard bond percolation process [85], each link in a network $G$ is retained with probability $\phi$ and removed with probability $1-\phi$, independently and uniformly at random. As $\phi$ varies, we track the size of the giant component—the largest connected subgraph—to study how connectivity responds to random removal of links. This process is relevant to various simple dynamics such as epidemics, information spreading, or behavior adoption, where deleted links represent blocked transmission paths. For example, in the case of epidemic spreading, under certain assumptions, the giant component corresponds precisely to the final outbreak size [94, 96]. As the disease becomes more infectious, $\phi$ increases, and the population undergoes a continuous phase transition from a contagion‐free to an endemic state. In more complex scenarios, link removal is not uniform and depends on node attributes. For example, disease transmission probabilities can differ based on age or vaccination status in a partially vaccinated population [97, 98, 99, 100, 101, 102, 103, 104, 14]. In such cases, percolation is no longer governed by a single parameter $\phi$, but by a vector of link-type-specific probabilities. In our colored network model, we remove links based on a vector $(\pi_{\mathrm{rr}},\pi_{\mathrm{rb}},\pi_{\mathrm{bb}})$, where $\pi_{kj}$ is the probability of removing a link between nodes of colors $k$ and $j$ [105]. In the large-network limit, this is equivalent to independently removing each $kj$-type link with probability $\pi_{kj}$. This is similar to the idea of neighbor-induced immunity percolation [106]. By varying these probabilities and observing how the giant component changes, we identify when the network undergoes a phase transition in connectivity.
Now, consider a network where red–blue and blue–blue links remain active with fixed probabilities, while the activation probability of red–red links varies from 0 to 1. In this setup, overall network connectivity is controlled by adjusting the presence of red–red links, $\pi_{\mathrm{rr}}$. Fig. 3(a–d) shows the size of the giant component for networks with the same Coleman index $h=0.3$, but different values of $h_{2}$ and $h_{6}$, which reflect varying levels of across- and within-group homophily.
When red-blue links are partially kept, $\pi_{\mathrm{rb}}>0$, panels (a) and (b), red and blue nodes can form a joint giant component altogether. Even with the same Coleman index $h$, networks with higher across-group homophily $h_{2}$ yield larger giant components at high $\pi_{\mathrm{rr}}$. Therefore, local mixing patterns affect not only the size of the giant component but also when the percolation transition happens, revealing structural effects that measures like the Coleman index miss. When red–blue links are absent, $\pi_{\mathrm{rb}}=0$, percolation can occur only through nodes of the same color, thus the size of the giant component is limited by the number of red or blue nodes. Different combinations of $h_{2}$ and $h_{6}$ modify the percolation threshold. In panel (d), where $\pi_{\mathrm{bb}}=0.2$, the giant component begins growing at the same threshold as in panel (c) ($\pi_{\mathrm{bb}}=0$), but from a higher baseline—indicating that percolation already occurs through blue–blue links, regardless of red-red links. Panel (a) shows that a large cluster emerges regardless of red-red links whenever a fraction of red-blue and blue-blue links are active. In panel (b), percolation thresholds are lower than in panel (c), due to the presence of across-group links. These results show that local mixing patterns affect not only the size of the giant component but also when the percolation transition happens, revealing structural effects that measures like the Coleman index miss.
@@ IV.1 Colored Percolation
Figure 4: Within- and across-group Homophily Influences Network Connectivity, Percolation Properties and Epidemic Thresholds in Non-Uniform Ways. Panels (a–c) present heatmaps of the critical percolation value with respect to $h=(h_{2}+h_{4})/2$ and $\Delta h=h_{2}-h_{4}$ for: (a) $\pi_{\mathrm{rr}}=\pi_{\mathrm{bb}}$, $\pi_{\mathrm{rb}}=0.1$; (b) $\pi_{\mathrm{rr}}=\pi_{\mathrm{bb}}$, $\pi_{\mathrm{rb}}=0.5$; (c) $\pi_{\mathrm{rr}}=0.1\pi_{\mathrm{bb}}$ $\pi_{\mathrm{rb}}=0.1,\alpha=0.2$. These panels illustrate how redistributing homophily configurations can either raise or lower the percolation threshold depending on the extent of across-group connectivity $\pi_{rb}$. Panel (d) compares the effect of redistributing homophily across small and large groups in networks with $h=0.5$ and average degree 2. Green regions indicate that emphasizing small-group homophily ($h_{4}=0.1$, $h_{2}=0.9$) leads to a higher percolation threshold than emphasizing large-group homophily ($h_{2}=0.1$, $h_{4}=0.9$). No percolation transition occurs in the top-right white region. Panels (e–g) explore the interaction between homophily and vaccination efficacy with respect to $h=(h_{2}+h_{4})/2$ and $\Delta h=h_{2}h_{4}$. The impact of homophily depends on both within-group efficacy ($f_{v}$) and cross-group efficacy ($f_{I}$): (e) $f_{I}=1$, $f_{v}=0.1$; (f) $f_{I}=0.1$, $f_{v}=0.5$; (g) $f_{I}=0.1$, $f_{v}=1$.
IV.2 Critical percolation value
@@
Analytical solutions for percolation on non-tree-like networks are famously intractable [107, 108]; yet, by isolating within-clique spreading from across-clique behavior, our framework breaks this impasse and provides a closed-form expression for the critical percolation threshold (see Section V.6). Using a multi-type branching process, we show that the percolation vector $\boldsymbol{\pi}=(\pi_{\mathrm{rr}},\pi_{\mathrm{rb}},\pi_{\mathrm{bb}})$ is critical when there exists a non-negative vector $\boldsymbol{v}$ such that
@@
We explore this condition in networks parametrized by within-group homophily $h_{4}$ and across-group homophily $h_{2}$. Fig. 4(a–b) shows heatmaps of the critical value $\pi_{\mathrm{rr}}=\pi_{\mathrm{bb}}$ for different values of $\pi_{\mathrm{rb}}$. When $\pi_{\mathrm{rb}}$ is low (Fig. 4a), increasing the global homophily expressed by the Coleman index $h$ reduces the critical threshold. When $\pi_{\mathrm{rb}}$ is high (Fig. 4b), increasing $h$ instead raises the threshold. Interestingly, these figures reveal how group homophilies affect connectivity, even when the Coleman index is fixed. We see that level curves for the critical value $\pi_{\mathrm{rr}}=\pi_{\mathrm{bb}}$ are not vertically aligned, meaning that for fixed $h$, distinct combinations of $h_{2}$ and $h_{4}$ yield different percolation thresholds. In particular, when $\pi_{\mathrm{rb}}$ is low, percolation is more easily achieved if $h_{4}>h_{2}$; vice versa, when $\pi_{\mathrm{rb}}$ is high, the percolation threshold is lower if $h_{2}>h_{4}$. The importance of local homophily is more prominent when $\pi_{rr}\neq\pi_{bb}$, as depicted in Fig. 4c. Here, the contour lines are roughly diagonal, indicating that the type of homophily is equally important to the amount of homophily for the percolation threshold. This means that the type of homophily can have a significant impact on the critical value of $\pi_{rr}$, which can range from 0.4 to 0.75 even when the Coleman homophily $h$ remains fixed.
@@
To illustrate the role of homophily within and across groups for the percolation threshold more systematically, we fix $h=0.5$ with $\alpha=0.5$ and compare two configurations: (i) weak across-group homophily ($h_{2}=0.1$) and strong within-group homophily ($h_{4}=0.9$), and (ii) the reverse. We compute the difference in critical values: $\Delta\pi_{\mathrm{rr}}^{*}=\pi_{\mathrm{rr}}^{*}(h_{4}=0.1)-\pi_{\mathrm{rr}}^{*}(h_{2}=0.1)$. Fig. 4d shows that this difference can be either positive or negative, depending on $\pi_{\mathrm{rb}}$ and $\pi_{\mathrm{bb}}$. Increasing the connectivity between groups (higher $\pi_{\mathrm{rb}}$ ) enhances the effectiveness of across-group homophily in maintaining network connectivity. In other words, when many cross-group links are present, networks with stronger across-group homophily achieve percolation at lower thresholds compared to networks with stronger within-group homophily. Thus, stronger homophily within or across groups does not universally improve connectivity. The difference in critical $\pi_{\mathrm{rr}}$ values between networks with the same Coleman index can be as large as 0.8, showing that global homophily alone is insufficient to predict percolation behavior.
IV.3 Epidemic Spread & Vaccination
@@
We demonstrate how multiscale homophily influences epidemic dynamics by applying our percolation framework to contact networks. Homophily affects disease transmission through vaccination behaviors and demographic mixing patterns alike. Age serves as a prime example: empirical contact matrices consistently show strong age-assortative mixing, with children primarily interacting with other children and adults exhibiting distinct interaction patterns depending on age and context [100, 101, 102, 103, 104]. Such structured contacts significantly reshape infection pathways, thereby affecting key epidemiological outcomes including herd immunity thresholds, expected epidemic sizes, and the effectiveness of interventions [14, 25, 26]. Our framework facilitates a detailed analysis of intervention strategies, providing a foundation for improved design and interpretation of public health policies [109]. To illustrate this, we analyze a vaccination scenario involving an imperfect vaccine, categorizing individuals as vaccinated (red) and unvaccinated (blue) nodes, to investigate how homophily shapes disease spread and influences the success of interventions.
@@
Fig. 4(e–g) shows that the impact of $f_{I}$ and $f_{v}$ on the critical percolation threshold is non-trivial. As in Fig. 4(a–c), the effect of homophily on the critical value can reverse. In Fig. 4e, increasing either homophily raises the critical threshold, while in Fig. 4g, it lowers it. This mirrors earlier results: for low $\pi_{\mathrm{rb}}$, increasing homophily decreases the critical value. Since $\pi_{\mathrm{rb}}=f_{I}\pi$, the values of $f_{I}$ in these panels explain the observed shift—$\pi_{\mathrm{rb}}=\pi$ in Fig. 4e and $0.1\pi$ in Fig. 4g. Fig. 4f shows that for certain combinations of $f_{I}$ and $f_{v}$, the critical value becomes nearly insensitive to either form of homophily.
@@
V Conclusion & Discussion
We have introduced a flexible framework for capturing homophily across multiple social scales, combining tightly knit subgroups with across-group interactions. By applying a maximum-entropy approach to maximal clique decomposition, our model accurately replicates homophily patterns in diverse real-world datasets, demonstrating how these trends remain hidden when homophily is considered as a single parameter. We further show that homophily may raise or lower percolation thresholds depending on the interaction between within- and across-group percolation, demonstrating that carefully tuned interventions (e.g., targeted vaccination or social distancing) can yield different outcomes. The connectivity insights provided by our framework remain applicable to any spreading phenomenon that can be effectively represented as a percolation process. Thus, our multiscale homophily decomposition provides a principled foundation for analyzing phenomena ranging from information diffusion to countering misinformation by managing link activation within and across groups.
@@ V Conclusion & Discussion
Nevertheless, the model can readily incorporate additional constraints on node degrees or hyperedge structures due to its underlying exponential-family formulation, allowing easy customization to empirical data. Furthermore, this flexibility extends naturally to scenarios where homophily parameters evolve over time presents a streamlined yet flexible approach for analyzing multiscale homophily, laying the groundwork for future studies that incorporateltiple attributes per node, capturing interactions between attributes that may result in non-additive effects, such as decreased homophily in one dimension influencing homophily in another [110]. Overall, this work presents a streamlined yet flexible approach for analyzing multiscale homophily, laying the groundwork for future studies that incorporate richer dynamics, degree heterogeneity, and higher-order motifs. This versatility enhances our understanding of how local assortative interactions influence global connectivity across social, biological, and technological networks, prompting further experimental and computational exploration into the broader implications of homophily-like biases in complex systems.
@@
Our decomposition of homophily reveals that what appears as modest or even absent assortativity at the network level often conceals a spectrum of structured, group-size-dependent mixing patterns. This observation challenges prevailing modeling assumptions and calls for a more elaborate understanding of how homophily operates in social systems and the implications for the various dynamics on them. By capturing this previously hidden structure in homophily patterns, our framework not only offers better alignment with empirical observations but also provides a principled model for investigating the behavior of dynamical processes on social networks.
@@ Acknowledgments
Funding: AKR acknowledges support from the Independent Research Fund Denmark (EliteForsk grant to Sune Lehmann), the Carlsberg Foundation (The Hope Project), and the Villum Foundation (NationScale Social networks). CS is supported by an NWO VENI grant 202.001. MK acknowledges the grant number 349366 from the Research Council of Finland. The simulations presented above were performed using computer resources within the Aalto University School of Science’s “Science-IT” project.
Use of AI tools: We utilized AI language tools to enhance grammar and readability. The authors reviewed and approved all content.
@@
Data and materials availability: All essential simulations and numerical computations are publicly available at [120]. The datasets used in this project were gathered from other research papers. Data collection methods, network details, and attribute distributions are provided in the original papers.

### v3

@@ III.2 Color Percolation
Figure 3: Within- and across-group Homophily Influences Network Connectivity, Percolation Properties and Epidemic Thresholds in Non-Uniform Ways. Panels (a–c) present heatmaps of the critical percolation value with respect to $h=(h_{2}+h_{4})/2$ and $\Delta h=h_{2}-h_{4}$ for: (a) $\pi_{\mathrm{rr}}=\pi_{\mathrm{bb}}$, $\pi_{\mathrm{rb}}=0.1$, and $\alpha_{2}=0.5$; (b) $\pi_{\mathrm{rr}}=\pi_{\mathrm{bb}}$, $\pi_{\mathrm{rb}}=0.5$, and $\alpha_{2}=0.5$; (c) $\pi_{\mathrm{rr}}=0.1\pi_{\mathrm{bb}}$ $\pi_{\mathrm{rb}}=0.1,\alpha_{2}=0.2$. These panels illustrate how redistributing homophily configurations can either raise or lower the percolation threshold depending on the extent of across-group connectivity $\pi_{rb}$. Panel (d) compares the effect of redistributing homophily across small and large groups in networks with $h=0.5$, average degree 2, and $\alpha_{2}=0.5$. Green regions indicate that emphasizing small-group homophily ($h_{4}=0.1$, $h_{2}=0.9$) leads to a higher percolation threshold than emphasizing large-group homophily ($h_{2}=0.1$, $h_{4}=0.9$). No percolation transition occurs in the top-right white region. Panels (e–g) explore the interaction between homophily and vaccination efficacy with respect to $h=(h_{2}+h_{4})/2$ and $\Delta h=h_{2}-h_{4}$. The impact of homophily depends on both within-group efficacy ($f_{v}$) and cross-group efficacy ($f_{I}$): (e) $f_{I}=1$, $f_{v}=0.1$; (f) $f_{I}=0.1$, $f_{v}=0.5$; (g) $f_{I}=0.1$, $f_{v}=1$.
In a standard bond percolation process [87], each link in a network $G$ is retained with probability $\phi$ and removed with probability $1-\phi$, independently and uniformly at random. As $\phi$ varies, we track the size of the giant component—the largest connected subgraph—to study how connectivity responds to random removal of links. This process is relevant to various simple dynamics such as epidemics, information spreading, or behavior adoption, where deleted links represent blocked transmission paths. For example, in the case of epidemic spreading, under certain assumptions, the giant component corresponds precisely to the final outbreak size [96, 98]. As the disease becomes more infectious, $\phi$ increases, and the population undergoes a continuous phase transition from a contagion‐free to an endemic state. In more complex scenarios, link removal is not uniform and depends on node attributes. For example, disease transmission probabilities can differ based on age or vaccination status in a partially vaccinated population [99, 100, 101, 102, 103, 104, 105, 106, 14]. In such cases, percolation is no longer governed by a single parameter $\phi$, but by a vector of link-type-specific probabilities. In our colored network model, we remove links based on a vector $(\pi_{\mathrm{rr}},\pi_{\mathrm{rb}},\pi_{\mathrm{bb}})$, where $\pi_{kj}$ is the probability of removing a link between nodes of colors $k$ and $j$ ($\pi_{kj}=\pi_{jk})$ [107]. In the large-network limit, this is equivalent to independently removing each $kj$-type link with probability $\pi_{kj}$. This is similar to the idea of neighbor-induced immunity percolation [108]. By varying these probabilities and observing how the giant component changes, we identify when the network undergoes a phase transition in connectivity.
III.3 Critical percolation value
@@
Analytical solutions for percolation on non-tree-like networks are famously intractable [109, 110]; yet, by isolating within-clique spreading from across-clique behavior, our framework breaks this impasse and provides a closed-form expression for the critical percolation threshold (see Section .7). Using a multi-type branching process, we show that the percolation vector $\boldsymbol{\pi}=(\pi_{\mathrm{rr}},\pi_{\mathrm{rb}},\pi_{\mathrm{bb}})$ is critical when there exists a non-negative vector $\boldsymbol{v}$ such that
@@
We explore this condition in networks parametrized by within-group homophily $h_{4}$ and across-group homophily $h_{2}$. Fig. 3(a,b) shows heatmaps of the critical value $\pi_{\mathrm{rr}}=\pi_{\mathrm{bb}}$ for different values of $\pi_{\mathrm{rb}}$. When $\pi_{\mathrm{rb}}$ is low (Fig. 3a), increasing the global homophily expressed by the Coleman index $h$ reduces the critical threshold. When $\pi_{\mathrm{rb}}$ is high (Fig. 3b), increasing $h$ instead raises the threshold. Interestingly, these figures reveal how group homophilies affect connectivity, even when the Coleman index is fixed. We see that level curves for the critical value $\pi_{\mathrm{rr}}=\pi_{\mathrm{bb}}$ are not vertically aligned, meaning that for fixed $h$, distinct combinations of $h_{2}$ and $h_{4}$ yield different percolation thresholds. In particular, when $\pi_{\mathrm{rb}}$ is low, percolation is more easily achieved if $h_{4}>h_{2}$; vice versa, when $\pi_{\mathrm{rb}}$ is high, the percolation threshold is lower if $h_{2}>h_{4}$. The importance of local homophily is more prominent when $\pi_{rr}\neq\pi_{bb}$, as depicted in Fig. 3c. Here, the contour lines are roughly diagonal, indicating that the type of homophily is equally important to the amount of homophily for the percolation threshold. This means that the type of homophily can have a significant impact on the critical value of $\pi_{rr}$, which can range from 0.4 to 0.75 even when the Coleman homophily $h$ remains fixed.
@@
To illustrate the role of homophily within and across groups for the percolation threshold more systematically, we fix $h=0.5$ with $\alpha_{2}=\alpha_{4}=0.5$ and compare two configurations: (i) weak across-group homophily ($h_{2}=0.1$) and strong within-group homophily ($h_{4}=0.9$), and (ii) the reverse. We compute the difference in critical values: $\Delta\pi_{\mathrm{rr}}^{*}$. Fig. 3d shows that this difference can be either positive or negative, depending on $\pi_{\mathrm{rb}}$ and $\pi_{\mathrm{bb}}$. Increasing the connectivity between groups (higher $\pi_{\mathrm{rb}}$ ) enhances the effectiveness of across-group homophily in maintaining network connectivity. In other words, when many cross-group links are present, networks with stronger across-group homophily achieve percolation at lower thresholds compared to networks with stronger within-group homophily. Thus, stronger homophily within or across groups does not universally improve connectivity. The difference in critical $\pi_{\mathrm{rr}}$ values between networks with the same Coleman index can be as large as 0.8, showing that global homophily alone is insufficient to predict percolation behavior.
III.4 Epidemic Spread & Vaccination
@@
We demonstrate how multiscale homophily influences epidemic dynamics by applying our percolation framework to contact networks. Homophily affects disease transmission through both vaccination behaviors and demographic mixing patterns. Age serves as a prime example: empirical contact matrices consistently show strong age-assortative mixing, with children primarily interacting with other children and adults exhibiting distinct interaction patterns depending on age and context [102, 103, 104, 105, 106]. Such structured contacts significantly reshape infection pathways, thereby affecting key epidemiological outcomes including herd immunity thresholds, expected epidemic sizes, and the effectiveness of interventions [14, 26, 27]. Our framework facilitates a detailed analysis of intervention strategies, providing a foundation for improved design and interpretation of public health policies [111]. To illustrate this, we analyze a vaccination scenario involving an imperfect vaccine, categorizing individuals as vaccinated (red) and unvaccinated (blue) nodes, to investigate how homophily shapes disease spread and influences the success of interventions.
@@
[…]
@@
IV Conclusion & Discussion
We introduced a generative maximum-entropy random-graph framework that captures homophily across multiple social scales with one parameter per group size. Sampling from the ensemble produces synthetic networks with prescribed multiscale homophily, and its exponential-family form enables maximum-likelihood inference of those parameters from data. Applied to diverse empirical datasets, the model integrates tightly knit groups with broader across-group interactions, accurately reproduces observed homophily patterns, and reveals structural insights that remain hidden when homophily is treated as a single global parameter. Subsequent simulation-based hypergraph work by Laber et al. [61] further underscores this scale dependence, showing that varying homophily across different interaction sizes creates pronounced disparities in groups’ timely access to information—an effect explained succinctly by our analytically tractable clique-layer approach.
Additionally, we demonstrated that homophily can either increase or decrease percolation thresholds, depending on the balance between within- and across-group connectivity. This result indicates that carefully designed interventions—such as targeted vaccination or selective social distancing—can produce significantly different outcomes. Since many spreading phenomena can be effectively modeled as percolation processes, our multiscale homophily decomposition provides a principled basis for analyzing and managing diffusion processes, from accelerating information spread to combating misinformation by selectively controlling interactions within and across groups.
@@ IV Conclusion & Discussion
The model’s exponential-family structure makes it easy to incorporate additional constraints—such as degree distributions, dynamic homophily parameters, and higher-order interactions—enabling adaptation to empirical data. It also supports extensions to multidimensional homophily, where traits interact in non-additive ways—especially when homophily in one dimension suppresses or enhances it in another [112]. This versatility provides a unified framework for studying how global connectivity emerges from local assortative preferences [113], offering a foundation for future work on richer dynamics, attribute interactions, and structural heterogeneity across social, biological, and technological networks.
@@
Our decomposition of homophily reveals that what appears as modest or even absent assortativity at the network level often conceals a spectrum of structured, group-size-dependent mixing patterns. This observation challenges prevailing modeling assumptions and calls for a more nuanced understanding of how homophily operates in social systems, as well as the implications for the various dynamics on them. By capturing this previously hidden structure in homophily patterns, our framework not only offers better alignment with empirical observations but also provides a principled model for investigating the behavior of dynamical processes on social networks.
@@ Acknowledgments
Funding: AKR acknowledges support from the Independent Research Fund Denmark (EliteForsk grant to Sune Lehmann), the Carlsberg Foundation (The Hope Project), and the Villum Foundation (NationScale Social networks). RM is supported by PNRR MUR project GAMING (PE0000013, CUP D13C24000430001). CS is supported by an NWO VENI grant 202.001. MK acknowledges the grant number 349366 from the Research Council of Finland. The simulations presented above were performed using computer resources within the Aalto University School of Science’s “Science-IT” project.
Data Availability: The datasets employed in this study were sourced from previously published research. Comprehensive descriptions of the data collection protocols, network specifications, and attribute distributions are available in the original publications. Readers are referred to the cited references for full details on each dataset.
Code Availability: Essential simulations and numerical computations are publicly available at [123].
Use of AI: We utilized AI language tools to enhance grammar and readability. The authors reviewed and approved all content.